UT 2021 Oct 26 - D/E Night



Night Info:

  • Split night between LBTI AO-onsky checkout, D/E tasks.
  • Night Center: 00:04
  • Handover from LBTI to LUCI-2 checkout: 23:59
  • Moon rise: 20:57
  • Sunset: 17:40
  • Twilight ends: 18:58
  • Twilight starts: 05:10
  • Sunrise 06:28


Night Plan:

  • LBTI on-sky checks until 23:59
    • thin clouds
    • seeing < 1.2"
  • LUCI-2 checkout starts 00:09
    • Rotator ZP - via astrometric masks - 0.5hr
      • thin clouds, seeing < 1.2"
      • Stone O (from UT 2021-10-07) - theta = -0.01 star#36 rejected)
      • Stone P sets 00:15 (30 deg)
      • Stone A sets 02:20 (30 deg) - worked best without any stars rejected for LUCI-1
      • Could test other LUCI MOS masks for extra datapoints or LS test
      • Files needed: ssh -X into obs1 as telescope; then ssh - X as tcs@tcs1
        • /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
        • Bounce GCS and PCS
    • AGW Transform - 1-1.5hrs
      • thin or variable clouds, seeing < 1.2"
      • Transform Collect
      • Potential lists:
        • Stone_A3_LUCI.list - sets 02:20
        • Stone_C1.list - RA ~04hr Dec -01 - sets twilight
        • Stone_D_new_Oct27.list - RA ~04:30hr Dec -00 - thru twilight
        • Stone E (-90,-135,90, 0 PAs) - RA 07hr Dec +00 - 01:30hr-twilight
  • LUCI-1 AO on-sky testing -1hr min - 2hr max
    • close the loop and let it run on-sky
    • Science-like AO Observation
      • Full AO
        • dither in-field, dither out-of-field
        • Rotator 20-80 degrees ONLY - rotator angle 45 best
        • ROTANGLE = POSANGLE - PARANGLE + (180 - ROTZPT for instrument)

        • Bin 1 R=12 point source to get rough Strehl - On-Axis
        • Bin 4 R=15 point source to get rough Stehl - On-Axis - POSTPONEDq
      • ESM
        • Bin 1 R<12
        • HCG 31 midnight-twilight
  • Additional Tasks:
    • LUCI (either) -0.25hr
      • Sub-frame readout test - crosstalk from bright stars outside the sub-array redout region
    • MODS
      • IT 8472 - GS at left edge of patrol field - 0.25hr
        • any conditions, but see below if photometric
        • Run sdss04+51_phot.txt in /home/lbto/Service_Observing/2021B_D_E/MODS
        • Takes 30-sec exposures at g+u(2x) and riz.(need star bright enough to WFS/guide on). We could make use of photometry since the field is in SDSS DR12.
    • LBCs
      • Active TMS - 3hr
        • Instructions
        • any conditions
        • start the night with a previous reference (the reference from 20211014), note the corrections indicated by dohybrid.
          • reference is tms?x_ref.dat_ and is in telescope@robs.mountain.lbto.org:~/TMS_logs/20211014/
          • on robs as user telescope, cp ~/TMS_logs/20211014/tms_?x_ref.dat /tmp
        • track a field through a large elevation change for ~2 hrs
          • Choose GD_279 or J031900+43. The scripts will dither 6x and change filters. You would need to repeat the OB about 4-5 times to make up 2 hours on the field.
          • visibility plot: 2021_oct26_2stds_vis.png
          • Scripts in: /home/lbto/Service_Observing/2021B_D_E/LBC/New_PhotFields/Scripts
          • Dual purpose: Please note whether it is photometric or not since these fields are from the set of new phot stds.
        • slew to a few fields at different elevations with TMS in active mode. Run FPIA and note the corrections indicated.
          • Use copointing fields for new phot standards.
          • Want to change elevation (see vis plot png or OT New_Phot_Fields_Master_OT_V1.1.xml).
            • visibility plot, 2021_oct_26_stds_vis.png.
  • Science Programs:
  • Poor Weather
    • LBTI:
      • up to 1.5" and thick clouds - 0.5 night AO checkout tasks
    • LBCs - TMS
    • MODS -

NightLog

LBTI

UT 00:37 - some issues with LBTI LMIRCam - com issue scrambles images on the GUIs (but actual images are OK) - if LBTI can't operate we will switch to LUCI-2 for checkout.

UT 00:38 - bringing up LUCI GUIs - inti all and darks

L1 0001-0002 - 2.51sec dark

L2 0011-0012 - 2.51 sec dark

UT 00:47 - Opening dome

UT 01:18 - starting up LBCs

-had to restart AO software for new pause/resume software

UT 01:27 - taking test LBC biases

UT 03:09 - closed for wind 22m/s gusts - right shutter door got stuck, eventually closed

UT 06:28 - LBTI wimping out and giving up telescope to the super-awesome LUCI tasks

LUCI-Bino

-reconfiguring for bino-LUCI

-particle counts up

UT 07:47 - particle counts down, but winds still high, humidity up to 75%

UT 08:01 - particle counts back up

UT 09:37 - now fogged over with 100% humidity. On the bright side winds are down just below 20 m/s.

UT 10:17 - water drops on all-sky

UT 10:37 - winds are back up over 20m/s.

UT 12:06 - turing off LBCs, closing out GUIs for LUCIs

...that's all folks...


-- BarryRothberg - 20 Oct 2021
Topic revision: r11 - 26 Oct 2021, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback